Express (I) i^(-4k) and (ii) i^(-(4k+1)), where i is the imaginary number and k is a positive constant,in the simplest form or in term of i.
Answer by Alan3354(69443)

Express (I) i^(-4k) and (ii) i^(-(4k+1)), where i is the imaginary number and k is a positive constant,in the simplest form or in term of i.
-------
i^(-4k) = 1/i^(4k) = 1/1 = 1
================================
i^(-(4k+1))
= 1/i^(4k+1)
= 1/(i^(4k)*i)
= 1/(1*i)
= 1/i
= i/i^2
= i/-1
= -i